# SuppSupplementary References

### v1

@@ 2.4 Exposure of a monotone curve in higher dimensions
The monotonicity of $\Gamma_{{\bf f}}$ is shown in Lemma LABEL:lemma1 in Supplementary LABEL:supp:proof. The proposition shows that any monotone curve in ${\mathbb{R}}^{k}$ can be exposed as a subset of the monotone set $\Gamma_{{\bf f}}$ induced by convex functions ${\bf f}$ in duality, and $\Gamma_{{\bf f}}$ is maximally monotone if and only if ${\bf f}$ is $c$-conjugate. This motivates us to use $H({\bf x};{\bf f})$ as the objective function. We note that Remark 2.1 can also be extended for ${\mathbb{R}}^{k}$. This is outlined in Supplementary LABEL:supp:extra.
@@ 3 Theoretical analysis
In this section, we first show the existence of our curve estimate found from the statistical learning problem and then investigate statistical error bounds for the estimate. The proof and the explicit forms of technical constants appear in Supplementary LABEL:supp:proof. To highlight the theoretical contributions of our work, we first review key literature in the field. We use the asymptotic notation $a_{n}=O(b_{n})$ (and $a_{n}=O_{p}(b_{n})$) to imply that $a_{n}/b_{n}$ is bounded above by a constant (and with high probability), as $n$ gets larger for the sequences $a_{n}$, $b_{n}$.
@@ 3.2 Statistical error analysis
In this section, we present our primary theoretical results on the statistical error bound of our estimate. We first outline the population model that generates the data. Consider a dimension $k\geq 2$, a maximally monotone curve $\Gamma\subset\mathbb{R}^{k}$, and a probability distribution $\rho\in\mathcal{P}(\mathbb{R}^{k})$ such that $\rho(\Gamma)=1$. Let $\varepsilon\in\mathcal{P}(\mathbb{R}^{k})$ denote the noise distribution with zero mean. Define ${\bf U}$ and ${\bf R}$ as independent random variables in $\mathbb{R}^{k}$, where ${\bf U}\sim\rho$ and ${\bf R}\sim\varepsilon$, thus ${\bf U}\in\Gamma$ and ${\mathbb{E}}[{\bf R}]=\textbf{0}$. The population model is then described by ${\bf X}={\bf U}+{\bf R}$ with ${\bf X}\sim\mu$, representing the population distribution. Let $\gamma(s)=\big{(}(\partial g_{i})^{-1}(s)\big{)}_{i\in[k]}$ parametrize $\Gamma$, where ${\bf f}=(f_{1},\dots,f_{k})$ is $c$-conjugate and exposes $\Gamma$, defined as in (LABEL:potential3) in Supplementary LABEL:supp:extra.
@@
Let ${\bf m}=(m_{0},m_{1},m_{2},m_{3})\in\mathbb{R}_{\geq 0}^{4}$ and ${\cal C}_{{\bf m}}$ be the class of admissible functions as in (6). Suppose $\Gamma\subset\mathbb{R}^{k}$ is maximally and strictly monotone, and its associated ${\bf f}=(f_{i})_{i\in[k]}$, defined in (LABEL:potential3), belongs to $({\cal C}_{{\bf m}})^{k}$. Assume ${\mathbb{E}}\|{\bf X}\|^{2}<\infty$. Given empirical data $\widehat{\mu}$, let $\widehat{{\bf f}}$ minimize (3) with ${\cal D}={\cal C}_{{\bf m}}$. Then for any data dimension $k\geq 2$ and the number of samples $n\in{\mathbb{N}}$, we have $E_{n}^{\textup{{emp}}}\leq C{\mathbb{E}}[\|{\bf R}\|^{2}]$ where $C>0$ depends only on $m_{2},\lambda$, and $k$. Moreover, if the components of the noise are mutually uncorrelated, i.e., ${\mathbb{E}}[R_{i}R_{j}]=0$ for every $i\neq j$, then $C$ depends only on $m_{2}$ and $\lambda$, and not on $k$.
We note that this dependence of the empirical error bound on noise size aligns with the argument in Hastie and Stuetzle (1989), which discusses how noise size amplifies the bias of the self-consistent curve, particularly when the curve is an arc of a circle.

### v2

@@ 2.4 Exposure of a monotone curve in higher dimensions
The monotonicity of $\Gamma_{{\bf f}}$ is shown in Lemma 8.1 in Supplementary 8. The proposition shows that any monotone curve in ${\mathbb{R}}^{k}$ can be exposed as a subset of the monotone set $\Gamma_{{\bf f}}$ induced by convex functions ${\bf f}$ in duality, and $\Gamma_{{\bf f}}$ is maximally monotone if and only if ${\bf f}$ is $c$-conjugate. This motivates us to use $H({\bf x};{\bf f})$ as the objective function. We note that Remark 2.1 can also be extended for ${\mathbb{R}}^{k}$. This is outlined in Supplementary 7.
@@ 3 Theoretical analysis
In this section, we first show the existence of our curve estimate found from the statistical learning problem and then investigate statistical error bounds for the estimate. The proof and the explicit forms of technical constants appear in Supplementary 8. To highlight the theoretical contributions of our work, we first review key literature in the field. We use the asymptotic notation $a_{n}=O(b_{n})$ (and $a_{n}=O_{p}(b_{n})$) to imply that $a_{n}/b_{n}$ is bounded above by a constant (and with high probability), as $n$ gets larger for the sequences $a_{n}$, $b_{n}$.
@@ 3.2 Statistical error analysis
In this section, we present our primary theoretical results on the statistical error bound of our estimate. We first outline the population model that generates the data. Consider a dimension $k\geq 2$, a maximally monotone curve $\Gamma\subset\mathbb{R}^{k}$, and a probability distribution $\rho\in\mathcal{P}(\mathbb{R}^{k})$ such that $\rho(\Gamma)=1$. Let $\varepsilon\in\mathcal{P}(\mathbb{R}^{k})$ denote the noise distribution with zero mean. Define ${\bf U}$ and ${\bf R}$ as independent random variables in $\mathbb{R}^{k}$, where ${\bf U}\sim\rho$ and ${\bf R}\sim\varepsilon$, thus ${\bf U}\in\Gamma$ and ${\mathbb{E}}[{\bf R}]=\textbf{0}$. The population model is then described by ${\bf X}={\bf U}+{\bf R}$ with ${\bf X}\sim\mu$, representing the population distribution. Let $\gamma(s)=\big{(}(\partial g_{i})^{-1}(s)\big{)}_{i\in[k]}$ parametrize $\Gamma$, where ${\bf f}=(f_{1},\dots,f_{k})$ is $c$-conjugate and exposes $\Gamma$, defined as in (9) in Supplementary 7.